# Multiple-policy Evaluation via Density Estimation

### v1

@@ 4.4 Main results
In this section, we provide our final sample complexity result of estimating the performance of all target policies based on the results we have in previous sections. We first give the main theorem and then provide the derivations.
@@
Implement Algorithm 2, then with probability at least $1-\delta$, for all target policies, we have that $|\hat{V}^{\pi^{k}}_{1}-V^{\pi^{k}}_{1}|\leq\epsilon$. And the total number of trajectories sampled is,
$n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$ (9)
Besides, the unknown true visitation distributions can be replaced by the coarse estimator to provide a concrete sample complexity,
$n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(\hat{d}_{h}^{\pi^{k}}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}\right)$
Now, we explain how the above result is derived. We first introduce a Median-of-Means (MoM) estimator and data splitting technique which can conveniently convert Lemma 4.7 to a version holds with high probability.
Lemma 4.9.
For a one-dimension value $\mu$, suppose we have a stochastic estimator $\hat{\mu}$ such that $\mathbb{E}[|\hat{\mu}-\mu|]\leq\frac{\epsilon}{4}$, then if we generate $N=O\left(\log(1/\delta)\right)$ i.i.d. estimators $\{\hat{\mu}_{1},\hat{\mu}_{2},\dots,\hat{\mu}_{N}\}$ and choose $\hat{\mu}_{MoM}=Median(\hat{\mu}_{1},\hat{\mu}_{2},\dots,\hat{\mu}_{N})$, we have with probability at least $1-\delta$,
$|\hat{\mu}_{MoM}-\mu|\leq\epsilon$
Proof.
See Appendix A.6. ∎
For step $h$, Algorithm 1 can output a solution $\hat{w}_{h}$ such that $\mathbb{E}[\ell^{\pi}_{h}(\hat{w}_{h})-\ell^{\pi}_{h}(w^{*}_{h})]\leq\frac{\epsilon^{2}}{32\xi H^{4}}$. We can apply Lemma 4.9 on our algorithm which means that we can run the algorithm for $N=O\left(\log(1/\delta)\right)$ times. Hence, we will get $N$ solutions $\{\hat{w}_{h,1},\hat{w}_{h,2},\dots,\hat{w}_{h,N}\}$. Set $\hat{w}_{h,MoM}$ as the solution such that $\ell^{\pi}_{h}(\hat{w}_{h,MoM})=Median(\ell^{\pi}_{h}(\hat{w}_{h,1}),\ell^{\pi}_{h}(\hat{w}_{h,2}),\dots,\ell^{\pi}_{h}(\hat{w}_{h,N}))$. Based on Lemma 4.9, we have that with probability at least $1-\delta$, it holds $\ell^{\pi}_{h}(\hat{w}_{h,MoM})-\ell^{\pi}_{h}(w^{*}_{h})\leq\frac{\epsilon^{2}}{32\xi H^{4}}$. With a little abuse of notation, we just denote $\hat{w}_{h,MoM}$ by $\hat{w}_{h}$ in the following content.
$\hat{V}^{\pi^{k}}_{1}=\frac{1}{n}\sum_{i=1}^{n}\sum_{h=1}^{H}\frac{\hat{w}^{\pi^{k}}_{h}(s_{h}^{i},a_{h}^{i})}{\hat{\mu}_{h}(s_{h}^{i},a_{h}^{i})}r_{h}(s_{h}^{i},a_{h}^{i})$ (10)
where $\{s^{i}_{h},a^{i}_{h}\}_{i=1}^{n}$ is sampled from $\tilde{\mu}_{h}$.
Lemma 4.10.
Proof.
Remember that in Section 4.1, we ignore those states and actions with low estimated visitation distribution for each target policy which induce at most $\frac{\epsilon}{2}$ error. Combined with Lemma 4.10, our estimator $\hat{V}^{\pi^{k}}_{1}$ finally achieves that with probability at least $1-\delta$, $|\hat{V}^{\pi^{k}}_{1}-V^{\pi^{k}}_{1}|\leq\epsilon,k\in[K]$.
Moreover, notice that,
$\max_{k\in[K]}\sum_{s,a}\frac{(\hat{d}^{\pi^{k}}_{h}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}\leq\max_{k\in[K]}\sum_{s,a}\frac{(\hat{d}^{\pi^{k}}_{h}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\leq\frac{25}{16}\sum_{s,a}\frac{(d^{\pi}_{h}(s,a))^{2}}{\mu^{*}_{h}(s,a)}$ (11)
@@
  Input: Accuracy $\epsilon$, confidence $\delta$, target policies $\{\pi^{k}\}_{k=1}^{K}$
  Roughly estimate visitation distributions of all target policies and get $\{\hat{d}^{\pi^{k}}\}_{k=1}^{K}$.
  Solve the approximate optimization problem (6) and get $\{\hat{\alpha}^{*}_{k}\}_{k=1}^{K}$.
  Implement Algorithm 1 and get $\{\hat{w}^{\pi^{k}}\}_{k=1}^{K}$.
@@
4.5 Discussion
In this section, we analyse our sample complexity and compare it with existing results. For off-policy evaluation, the CR-lower bound proposed by Jiang and Li (2016) (Theorem 3) shows that there exists a MDP such that the variance of any unbiased estimator is lower bounded by
$\sum_{h=1}^{H}\mathbb{E}_{\mu}\left[\left(\frac{d^{\pi}_{h}(s_{h},a_{h})}{\mu_{h}(s_{h},a_{h})}\right)^{2}\mathbb{V}[V^{\pi}_{h}(s_{h})]\right],$
where $\pi$ is the policy to evaluate and $\mu$ is the sampling distribution generated by some behavior policy. If we simply bound the variance of the value function by $H^{2}$ due to the boundness of rewards and omit the dependency on $H$ which we will discuss later, our sample complexity matches the lower bound for multiple-policy evaluation since our sampling distribution is optimal.
@@ 4.5 Discussion
One interesting proposition from our sample complexity result is that in the case where all target policies are identical, i.e. $d^{\pi^{1}}=d^{\pi^{2}}=\cdots=d^{\pi^{K}}=d$. Then, the optimal sampling distribution is $\mu^{*}=d$, hence, the sample complexity is $\frac{H^{5}}{\epsilon^{2}}$ which has no dependency on $S$ or $A$.
@@
We can also derive a non-instance dependent sample complexity based on our results. Let the sampling distribution $\mu^{\prime}_{h}$ be $\frac{1}{SA}\sum_{s,a}d^{\pi_{s,a}}_{h}$, where $\pi_{s,a}=\arg\max_{k\in[K]}d^{\pi^{k}}_{h}(s,a)$. Since $\mu^{*}_{h}$ is the optimal solution and $\mu^{\prime}_{h}$ is a feasible solution, we have our sample complexity (9) is bounded,
$n\leq\tilde{O}\left(\frac{H^{5}SA}{\epsilon^{2}}\right).$
Now, we compare our result with the one achieved by Dann et al. (2023). One big problem of the result by Dann et al. (2023) is the existence of unfavorable $\frac{1}{d^{max}(s)}$ which can induce an undesired dependency on $K$. We next illustrate it by an example. Consider such a MDP with two layers, a single initial state $s_{1,1}$ in the first layer and two terminal states in the second layer $s_{2,1},s_{2,2}$. The transition function is same for all actions, i.e. $P(s_{2,1}|s_{1,1},a)=p$ and $p$ is sufficiently small. The agents only receive rewards at state $s_{2,1}$ no matter what actions they take. Hence, to evaluate the performance of some policy under this MDP, it is enough to only consider the second layer.
Now, suppose we have $K$ target policies to evaluate such that they take different actions at state $s_{1,1}$ and take the same action at any state in the second layer. Since the transition function at state $s_{1,1}$ is same for any action, the visitation distribution at state $s_{2,1}$ of all target policies is identical.
Since $p$ is sufficiently small, we have that the probability of $s_{2,1}$ being reached is $\mathbb{P}[s_{2,1}\in\mathcal{K}^{2}]=1-(1-p)^{K}\approx pK$. Invoke the result (1) of Dann et al. (2023), the sample complexity in this case has a dependency on $K$. At the same time, our result can show the sample complexity without dependency on $K$. However, it is unclear whether our result is better in all cases.
@@
Finally, comparing with the naive uniform sampling strategy over target policies, our method is clearly at advantage since our sampling distribution is optimal within all possible combinations of the target policies which include the naive uniform one.
@@ 5 Conclusion and Future Work
In this work, we consider the problem of multi-policy evaluation. And we propose an algorithm $\mathrm{CAESAR}$ based on computing an approximate optimal offline sampling distribution and using the data sampled from it to perform the simultaneous estimation of the policy values. $\mathrm{CAESAR}$ achieves that with number of trajectories $n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$ and probability at least $1-\delta$, we can evaluate the performance of all target policies up to $\epsilon$ error. The algorithm consists of three techniques. First, we obtain a coarse distribution estimator at the cost of lower-order sample complexity. Second, based on the coarse distribution estimator, we show an achievable optimal sampling distribution by solving an convex optimization problem. Last, we propose a novel step-wise loss function for finite-horizon MDPs. By minimizing the loss function step to step, we are able to get the importance weighting ratio and a non-asymptotic sample complexity is available due to the smoothness and strongly-convexity of the loss function.
Beyond the result of this work, there are still some open questions of interest. First, our sample complexity has a dependency on $H^{4}$ which is induced by the error propagation in the estimation of the importance weighting ratios. Specifically, the error of minimizing the loss function at early steps, e.g $h=1$ will propagate to later steps e.g $h=H$. We conjecture a dependency on $H^{2}$ is possible by considering a comprehensive loss function includes the whole horizon instead of step-wise loss functions which require step by step optimization. Besides, considering a reward-dependent sample complexity is also an interesting direction. For example, consider a MDP with sparse rewards where only one state-action has non-zero reward, then a better sample complexity may be possible by just focusing on state-action pairs with non-zero rewards. Another future direction is to apply the coarse distribution estimator on more scenarios. In our work, the coarse distribution estimator plays an important role throughout the algorithm. And we believe this type of estimator has potentiality in other different settings and tasks.
Acknowledgements.
@@ Appendix A Proof of theorems and lemmas in Section 4
Our results relies on the following variant of Bernstein inequality for martingales, or Freedman’s inequality Freedman [1975], as stated in e.g., Agarwal et al. [2014], Beygelzimer et al. [2011].
@@
Consider the martingale difference sequence $X_{t}=Z_{t}-\mathbb{E}_{t}[Z_{t}]$. Notice that $|X_{t}|\leq\tilde{B}$. Using the inequality of Lemma A.2 we obtain for all $\eta\in(0,1/B^{2})$.
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}2\eta B^{2}\sum_{\ell=1}^{t}\mathbb{E}_{\ell}[Z_{\ell}]+\frac{C\log(t/\delta^{\prime})}{\eta}$
@@
$\leq 2\eta B^{2}\sum_{\ell=1}^{t}\mathbb{E}_{\ell}[Z_{\ell}]+\frac{C\log(t/\delta^{\prime})}{\eta}$
[…]
@@
with probability at least $1-\delta^{\prime}$. Combining Equations 14 and 15 and using a union bound yields the desired result.
Proposition A.3 can be used to show,
@@
$(1-\beta)p-\epsilon/8\leq\widehat{p}_{t}\leq(1+\beta)p+\epsilon/8$
@@
$87p/8-\epsilon/8\leq\widehat{p}_{t}\leq 9p/8+\epsilon/$ 8
@@
$\widehat{p}_{t}-p\leq p/8+\epsilon/8.$
and therefore $|\widehat{p}_{t}-p|\leq p/8+\epsilon/8\leq 2\max(p/8,\epsilon/8)=\max(p/4,\epsilon/4)$.
A.2 Proof of the optimal sampling distribution (4)

### v2

@@ 4.4 Main results
We are now ready to present our main sample complexity result for multiple-policy evaluation, building on the results from previous sections. First, we introduce a Median-of-Means (MoM) estimator [20], formalized in the following lemma, and a data splitting technique that together can be used to convert (8) into a high-probability result (see Appendix A.7).
Lemma 4.7.
Let $\mu\in\mathbb{R}$ and suppose we have a stochastic estimator $\hat{\mu}$ such that $\mathbb{E}[|\hat{\mu}-\mu|]\leq\frac{\epsilon}{4}$. Then, if we generate $N=O\left(\log(1/\delta)\right)$ i.i.d. estimators $\{\hat{\mu}_{1},\hat{\mu}_{2},\dots,\hat{\mu}_{N}\}$ and choose $\hat{\mu}_{MoM}=\text{Median}(\hat{\mu}_{1},\hat{\mu}_{2},\dots,\hat{\mu}_{N})$, we have with probability at least $1-\delta$,
$|\hat{\mu}_{MoM}-\mu|\leq\epsilon.$
With the importance density estimator $\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}$, we can estimate the performance of policy $\pi^{k}$,
$\hat{V}^{\pi^{k}}_{1}=\frac{1}{n}\sum_{i=1}^{n}\sum_{h=1}^{H}\frac{\hat{w}^{\pi^{k}}_{h}(s_{h}^{i},a_{h}^{i})}{\hat{\mu}_{h}(s_{h}^{i},a_{h}^{i})}r_{h}(s_{h}^{i},a_{h}^{i}).$ (9)
where $\{s^{i}_{h},a^{i}_{h}\}_{i=1}^{n}$ is sampled from $\tilde{\mu}_{h}$.
We present our main result in the following theorem and leave its detailed derivation to Appendix A.7.
@@
Implement Algorithm 2. Then, with probability at least $1-\delta$, for all target policies, we have that $|\hat{V}^{\pi^{k}}_{1}-V^{\pi^{k}}_{1}|\leq\epsilon$. And the total number of trajectories sampled is,
$n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right).$ (10)
Besides, the unknown true visitation distributions can be replaced by the coarse estimator to provide a concrete sample complexity.
@@
  Input: Accuracy $\epsilon$, confidence $\delta$, target policies $\{\pi^{k}\}_{k=1}^{K}$.
  Coarsely estimate visitation distributions of all deterministic policies and get $\{\hat{d}^{\pi}:\pi\in\Pi_{det}\}$.
  Solve the approximate optimization problem (6) and obtain $\{\hat{\alpha}^{*}_{\pi}:\pi\in\Pi_{det}\}$.
  Implement Algorithm 1 with data splitting and obtain MoM estimators $\{\hat{w}^{\pi^{k}}\}_{k=1}^{K}$.
@@
5 Discussion
In this section, we analyze our sample complexity, comparing it with existing results and offering several noteworthy findings.
5.1 Lower bound and some special cases
For off-policy evaluation, the CR-lower bound proposed by [16] (Theorem 3) demonstrates that there exists an MDP such that the variance of any unbiased estimator is lower bounded by $\sum_{h=1}^{H}\mathbb{E}_{\mu}\left[\left(\frac{d^{\pi}_{h}(s_{h},a_{h})}{\mu_{h}(s_{h},a_{h})}\right)^{2}\mathbb{V}[V^{\pi}_{h}(s_{h})]\right]$, where $\pi$ is the policy to evaluate and $\mu$ is the sampling distribution. Applying this result to multiple-policy evaluation problem gives us the lower bound $\min_{\mu}\max_{k\in[K]}\sum_{h=1}^{H}\mathbb{E}_{\mu}\left[\left(\frac{d^{\pi^{k}}_{h}(s_{h},a_{h})}{\mu_{h}(s_{h},a_{h})}\right)^{2}\mathbb{V}[V^{\pi^{k}}_{h}(s_{h})]\right]$. From the variance-unaware perspective, where the variance of the value function is simply bounded by $H^{2}$, our sample complexity matches this lower bound since our sampling distribution is optimal (up to the dependency on $H$). We believe that a more refined variance-dependent result is achievable and leave it to future works.
@@ 5.1 Lower bound and some special cases
Next, we analyse our sample complexity based on some special cases which offers us some interesting results. First, in the scenario where all target policies are identical, i.e. $d^{\pi^{1}}=d^{\pi^{2}}=\cdots=d^{\pi^{K}}=d$. The optimal sampling distribution is $\mu^{*}=d$, hence, our sample complexity becomes $\tilde{O}(\frac{H^{5}}{\epsilon^{2}})$ which has no dependency on $S$ or $A$.
@@
We can derive an instance-independent sample complexity based on our results. Let the sampling distribution $\mu^{\prime}_{h}$ be $\frac{1}{SA}\sum_{s,a}d^{\pi_{s,a}}_{h}$, where $\pi_{s,a}=\arg\max_{k\in[K]}d^{\pi^{k}}_{h}(s,a)$. Since $\mu^{*}_{h}$ is the optimal solution and $\mu^{\prime}_{h}$ is a feasible solution, we have our sample complexity (10) is bounded by $\tilde{O}\left(\frac{H^{5}SA}{\epsilon^{2}}\right)$.
5.2 Comparison with existing results
@@
First, compared to the naive uniform sampling strategy over target policies as described in (2), our method has a clear advantage. Our sampling distribution is optimal among all possible combinations of the target policies, including the naive uniform strategy.
Next, we compare our result with the one achieved by Dann et al. [5] as described in (1). A significant issue with the result by Dann et al. [5] is the presence of the unfavorable $\frac{1}{d^{max}(s)}$, which can induce an undesirable dependency on $K$.
To illustrate this, consider an example of an MDP with two layers: a single initial state $s_{1,1}$ in the first layer and two terminal states in the second layer $s_{2,1},s_{2,2}$. The transition function is the same for all actions, i.e., $P(s_{2,1}|s_{1,1},a)=p$ and $p$ is sufficiently small. Agents only receive rewards at state $s_{2,1}$, regardless of the actions they take. Hence, to evaluate the performance of a policy under this MDP, it is sufficient to consider only the second layer. Now, suppose we have $K$ target policies to evaluate, where each policy takes different actions at state $s_{1,1}$ but the same action at any state in the second layer. Since the transition function at state $s_{1,1}$ is the same for any action, the visitation distribution at state $s_{2,1}$ of all target policies is identical. Given that $p$ is sufficiently small, the probability of reaching $s_{2,1}$ is $\mathbb{P}[s_{2,1}\in\mathcal{K}^{2}]=1-(1-p)^{K}\approx pK$. According to the result (1) by Dann et al. [5], the sample complexity in this scenario depends on $K$. In contrast, our result demonstrates sample complexity without dependency on $K$. However, it remains unclear whether our result is universally better in all cases (omit the dependency on $H$).
5.3 Near-optimal policy identification
Besides policy evaluation, $\mathrm{CAESAR}$ can also be applied to identify a near-optimal policy. Fixing the high-probability factor, we denote the sample complexity of $\mathrm{CAESAR}$ by $\tilde{O}(\frac{\Theta(\Pi)}{\gamma^{2}})$, where $\Pi$ is the set of policies to be evaluated and $\gamma$ is the estimation error. We provide a simple algorithm based on $\mathrm{CAESAR}$ in Appendix D that achieves an instance-dependent sample complexity $\tilde{O}(\max_{\gamma\geq\epsilon}\frac{\Theta(\Pi_{\gamma})}{\gamma^{2}})$ to identify a $\epsilon-$optimal policy, where $\Pi_{\gamma}=\{\pi:V^{*}_{1}-V^{\pi}_{1}\leq 8\gamma\}$. This result is interesting as it offers a different perspective beyond the existing gap-dependent results [22, 7]. Furthermore, this result can be easily extended to the multi-reward setting. We leave the detailed discussion to Appendix D.
@@ 6 Conclusion and Future Work
In this work, we consider the problem of multi-policy evaluation. We propose an algorithm, $\mathrm{CAESAR}$ , based on computing an approximate optimal offline sampling distribution and using the data sampled from it to perform the simultaneous estimation of the policy values. $\mathrm{CAESAR}$ uses $n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$ trajectories and with probability at least $1-\delta$ we can evaluate the performance of all target policies up to an $\epsilon$ error. The algorithm consists of three techniques. First, we obtain a coarse distribution estimator at the cost of lower-order sample complexity. Second, based on the coarse distribution estimator, we show an achievable optimal sampling distribution by solving a convex optimization problem. Last, we propose a novel step-wise loss function for finite-horizon MDPs. By minimizing the loss function step to step, we are able to get the importance weighting ratio and a non-asymptotic sample complexity is available due to the smoothness and strong-convexity of the loss function.
Beyond the results of this work, there are still some open questions of interest. First, our sample complexity has a dependency on $H^{4}$ which is induced by the error propagation in the estimation of the importance weighting ratios. Specifically, the error of minimizing the loss function at early steps, e.g., $h=1$ will propagate to later steps e.g., $h=H$. We conjecture a dependency on $H^{2}$ is possible by considering a comprehensive loss function which includes the entire horizon instead of step-wise loss functions which require step by step optimization. Second, as discussed before, we believe that a variance-aware sample complexity is possible through a more careful analysis. Besides, considering a reward-dependent sample complexity is also an interesting direction. For example, in an MDP with sparse rewards where only one state-action has non-zero reward; a better sample complexity than $\mathrm{CAESAR}$ ’s may be possible by focusing on state-action pairs with non-zero rewards. Finally, we are interested to see what other uses the research community may find for coarse distribution estimation. In our work, the coarse distribution estimator plays an important role throughout the algorithm. We believe this type of estimator can be of independent interest.
Acknowledgements
@@ Appendix A Proof of theorems and lemmas in Section 4
Our results relies on the following variant of Bernstein inequality for martingales, or Freedman’s inequality [12], as stated in e.g., [1, 3].
@@
Consider the martingale difference sequence $X_{t}=Z_{t}-\mathbb{E}_{t}[Z_{t}]$. Notice that $|X_{t}|\leq\tilde{B}$. Using the inequality of Lemma A.2 we obtain that for all $\eta\in(0,1/B^{2})$
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}2\eta B^{2}\sum_{\ell=1}^{t}\mathbb{E}_{\ell}[Z_{\ell}]+\frac{C\log(t/\delta^{\prime})}{\eta},$
@@
$\leq 2\eta B^{2}\sum_{\ell=1}^{t}\mathbb{E}_{\ell}[Z_{\ell}]+\frac{C\log(t/\delta^{\prime})}{\eta}.$
Setting $\eta=\frac{\beta}{2B^{2}}$ and substituting $\sum_{\ell=1}^{t}X_{\ell}^{\prime}=\sum_{\ell=1}^{t}\mathbb{E}[Z_{\ell}]-Z_{\ell}$ we have,
@@
with probability at least $1-\delta^{\prime}$. Combining Equations 13 and 14 and using a union bound yields the desired result.
@@
$(1-\beta)p-\epsilon/8\leq\widehat{p}_{t}\leq(1+\beta)p+\epsilon/8.$
@@
$7p/8-\epsilon/8\leq\widehat{p}_{t}\leq 9p/8+\epsilon/8,$
@@
$\widehat{p}_{t}-p\leq p/8+\epsilon/8,$
which implies $|\widehat{p}_{t}-p|\leq p/8+\epsilon/8\leq 2\max(p/8,\epsilon/8)=\max(p/4,\epsilon/4)$.
A.2 Derivation of the optimal sampling distribution (4)